OpenStudy (anonymous):

calculate the molecular mass of the solute when 60.5 grams of a non-ionizing solute in 450 grams of water raises the boiling point to 101 degrees celsius

OpenStudy (aaronq):

do you know what equation to use?

OpenStudy (anonymous):

i think its ΔT b.p.=(0.515)(m)(i)

OpenStudy (aaronq):

yep, thats exactly it. all you need to do is a substitution and then an algebraic rearrangement. molality=moles of solute/mass of solvent in Kg for moles, n, sub in: moles =mass/molar mass (i.e. n=m/M)

OpenStudy (anonymous):

im confused on how you do that

OpenStudy (aaronq):

molality = moles/kg of solvent sub in n=m/M molality=(m/M)/kg of solvent

OpenStudy (aaronq):

you have to rearrange to solve for M, molecular mass

OpenStudy (anonymous):

ya im still lost....

OpenStudy (aaronq):

lol \[\Delta T = i T _{b}m \rightarrow \Delta T = i T _{b}\frac{ \frac{ m }{ M } }{ kg _{solvent} }\]

OpenStudy (aaronq):

remember m in the second equation means mass, not molality